A population with a size of 4 has a mean of 5.41 and a variance of 6.8. The sampling distribution has a sample size of 3 and all possible samples are drawn without replacements. Find the mean, variance, and standard deviation.


1
Expert's answer
2021-06-28T16:36:17-0400

(a). The mean of sampling distribution is:

"\mu \:_{\overline{x}}=\frac{sum\:of\:all\:samples}{number\:of\:samples}=\frac{6}{3}=2"


(b). The variance of sampling distribution is:

The variance "\sigma ^2_{\overline{x}}" of the sampling distribution means is obtained by subtracting the mean 2 from each number, squaring the result, adding all 3 numbers obtained, and dividing by 3. The final result is:

"\sigma ^2_{\overline{x}}=\frac{2}{3}=0.6667"


(c). The standard deviation of sampling distribution is:

Standard deviation is the square root of variance obtained in b

"\sigma _{\overline{x}}=\sqrt{\sigma ^2_{\overline{x}}}"

"\sigma \:_{\overline{x}}=\sqrt{0.6667}=0.8165"